OpenStudy (anonymous):

thanks, now I have a second question. please factorize: (x2 - 2xy+y2) - z2

OpenStudy (anonymous):

= (x-y)^2 - z^2 which could be further factored to (x-y-z)(x-y+z)

OpenStudy (anonymous):

= (x - y)² - z² = [(x-y) + z][(x-y) - z] = (x - y + z) (x - y - z)

OpenStudy (anonymous):

johny, but my question is (x2-2xy+y2)-z2

OpenStudy (anonymous):

thanks, got it.

OpenStudy (anonymous):

then click GOOD ANSWER button pls...☺